If the price of coffee was 50 cents a pound 10 years ago, and the current price of that product is $4.25 a pound, what percent increase in price does this represent?
Answer by amarjeeth123(569)

price of coffee was 50 cents($0.50) a pound 10 years ago
Current price=$4.25 a pound
Increase=4.25-0.50=$3.75
Percent increase=3.75/0.5*100=750%